OpenStudy (anonymous):

The length of a rectangle is 8 inches and the width is x inches. What is the perimeter, in inches?

OpenStudy (igreen):

Well, you can't find the Perimeter of the rectangle, because you have 2 variables. The formula for the Perimeter of a rectangle is: \(P = 2l + 2w\) Plug in what we know: \(P = 2(8) + 2(x)\) Simplify: \(P = 2x + 16\) So that's basically your answer.. @JulianneSantos

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Resposne'. @JulianneSantos